# TacMan-Turbo: Proactive Tactile Control for Robust and Efficient Articulated Object Manipulation

### v2

@@ Abstract
[…]
keywords:
Contact control, contact sensing, manipulation of articulated objects, and efficient manipulation.
11footnotetext: Zihang Zhao and Zhenghao Qi contributed equally to this work.
Figure 1: Robust and efficient tactile-informed manipulation of articulated objects. TacMan-Turbo enables robots to achieve both effective and efficient manipulation of articulated objects through a novel tactile-based control framework.Unlike previous approaches, TacMan-Turbo leverages past tactile experience to extract local kinematic information, allowing the controller to predict optimal interaction positions and make proactive adjustments—all without requiring pre-provided object kinematics. This capability significantly enhances manipulation efficiency while maintaining robustness across diverse scenarios. Panels (a-c) demonstrate TacMan-Turbo’s versatility across various joint mechanisms: (a) manipulating a prismatic-joint drawer in just 10 seconds, (b) operating a revolute-joint microwave oven in 24 seconds, and (c) controlling a complex vise with a helical handle path in 24 seconds. Panel (d) quantifies these improvements, showing that while maintaining 100% effectiveness, our method achieves approximately $5\times, 11$$1\times$, and $18\times$ higher time efficiency than the recent tactile-informed method, Tac-Man [73], as measured by average manipulation speeds. These order-of-magnitude improvements in efficiency without compromising reliability effectively resolve the longstanding trade-off between effectiveness and efficiency in articulated object manipulation.
1 Introduction
@@
Adept manipulation of articulated objects, from doors and drawers to various household items, constitutes one of the basic capabilities that robots must master to function effectively in human-centric environments [54]. Success in this domain hinges on two critical objectives: effectiveness—the reliable manipulation of objects despite uncertain internal structures, and efficiency—the execution of tasks with minimal redundant movements and smooth trajectories [24, 23].
@@
Current approaches struggle to achieve both objectives simultaneously. Methods that rely on pre-provided kinematic information demonstrate efficiency in controlled scenarios but often fail when encountering discrepancies between assumed and actual object kinematics. Conversely, while the recent tactile-informed approach by Zhao et al[73] successfully achieves robust manipulation despite kinematic uncertainty, it sacrifices efficiency—requiring several minutes to manipulate a single object due to its cautious, step-by-step exploration-compensation mechanism. This fundamental trade-off between effectiveness and efficiency remains unresolved in articulated object manipulation, as further detailed in Section 2.1.
To address this challenge, we propose TacMan-Turbo, a novel proactive control framework that fundamentally reimagines how tactile feedback can guide manipulation. While acknowledging Zhao et al.’s insight [73] in using robot-object contact deviations as state representations, TacMan-Turbo interprets these tactile-informed deviations in a different manner—not just as position errors requiring compensation, but as rich sources of information indicating the object’s local kinematics. By systematically collecting sequences of these observations, TacMan-Turbo enables accurate prediction of subsequent interaction positions with local optimality. This predictive capability allows the controller to make proactive adjustments rather than relying solely on reactive corrections, thereby significantly improving manipulation efficiency while maintaining robustness. As illustrated in Fig. 1, this novel approach preserves the effectiveness of contact-based manipulation while enabling substantially more efficient execution.
@@
We validate TacMan-Turbo through comprehensive evaluations in both simulation and real-world settings. Our test scenarios systematically cover articulated objects with diverse joint types—from basic prismatic and revolute joints to complex mechanisms involving concurrent translational and rotational movements. Across $\boldsymbol{200}$ simulated articulated objects and multiple real-world experiments, TacMan-Turbo achieves the same perfect manipulation success rate as the previous tactile-informed approach while significantly improving all efficiency metrics: time efficiency, action efficiency, and trajectory smoothness (all comparisons yielding p-values $\boldsymbol{0.0001}$). These results confirm that TacMan-Turbo successfully mitigates the long-standing effectiveness-efficiency trade-off in articulated object manipulation.
This paper makes two major contributions:
  • We introduce TacMan-Turbo, a novel proactive tactile control framework that effectively resolves the trade-off between efficiency and effectiveness in articulated object manipulation. While traditional approaches treat contact deviations solely as error signals requiring compensation, TacMan-Turbo extracts valuable kinematic information from these deviations to enable proactive adjustments through robust tactile-informed predictive control. This paradigm shift significantly enhances manipulation efficiency while maintaining effectiveness and eliminating reliance on pre-provided object kinematics—data that is often imprecise or unavailable in dynamic, human-centric environments.
  • Through extensive evaluation across $\boldsymbol{200}$ diverse articulated objects in simulation and multiple real-world experiments, we demonstrate TacMan-Turbo’s ability to achieve $100\%$ manipulation success rate while significantly improving all efficiency metrics—time consumption, action efficiency, and trajectory smoothness—regardless of joint configurations or kinematic complexities. Our comprehensive real-world validation confirms that this enhanced efficiency occurs without compromising manipulation reliability, representing a meaningful advancement toward practical robot deployment in human environments.
@@
The remainder of this paper is organized as follows: Section 2 positions our work within the broader context of robotic manipulation research. Section 3 presents the theoretical framework and implementation details of TacMan-Turbo. Sections 4 and 5 evaluate TacMan-Turbo through simulation studies and real-world experiments, respectively. Section 6 analyzes TacMan-Turbo’s performance characteristics and discusses future research directions. Finally, Section 7 summarizes our findings and contributions.
2 Related Work
@@
This section contextualizes TacMan-Turbo’s contributions within the broader landscape of robotic research. We begin by assessing current approaches to articulated object manipulation, establishing the fundamental trade-off between effectiveness and efficiency that motivates our work (Section 2.1). In Section 2.2, we explore contact modeling techniques, highlighting how the proposed method builds upon and advances previous work to address the effectiveness-efficiency trade-off. Next, Section 2.3 examines contact sensing technologies, categorizing them into force-based and geometry-based approaches while explaining their compatibility with our method. Finally, Section 2.4 reviews contact control strategies, tracing their evolution from early reactive feedback systems to our proposed proactive tactile control frameworks.
2.1 Articulated Object Manipulation
@@
To overcome this dependency on manually provided information, subsequent research has focused on the autonomous acquisition of object kinematics. A prominent line of work leverages visual perception to infer kinematic models through geometric reasoning and motion analysis [20, 29, 68, 40, 10, 34, 44]. However, visual information alone often proves insufficient, as visually similar objects may possess fundamentally different internal mechanisms [75, 68, 73]. This inherent limitation has led researchers to integrate additional sensing modalities to resolve ambiguity. By combining multi-frame analysis with wrist-mounted force/torque sensors, more accurate parameter estimation of kinematic models becomes possible through active interaction [24, 19, 23, 39, 36, 31]. Nevertheless, these methods remain constrained by their underlying assumption that objects conform to predefined joint types (e.g., revolute, prismatic), limiting their effectiveness when encountering complex or unconventional mechanisms.
@@
Reinforcement learning offers an alternative paradigm that attempts to overcome the data collection bottleneck by developing manipulation strategies through repeated interactions in simulated environments [53, 62, 7, 15]. This approach has been significantly advanced by the creation of large-scale datasets of articulated objects [41, 60, 33, 16], enabling training across diverse object categories. However, the inevitable sim-to-real gap often leads to degraded performance when these methods are deployed in real-world settings, particularly when encountering previously unseen object variations.
The limitations in both model-based and learning-based approaches recently led to the development of a tactile-informed manipulation framework. Zhao et al[73] introduced a novel approach that achieves reliable manipulation by formulating it as a control problem that compensates for robot-object contact deviations without requiring kinematic priors. While this method demonstrates remarkable effectiveness in handling unknown articulated objects, its step-by-step exploration-compensation process results in significant time overhead that limits practical deployment. Our work, TacMan-Turbo, advances this foundation by extracting structural information from tactile feedback to enable predictive control. This proactive approach significantly improves manipulation efficiency while maintaining the same level of effectiveness—effectively addressing the long-standing trade-off between these two critical objectives.
2.2 Contact Modeling
Contact modeling has emerged as a fundamental approach in robotics research, providing essential information about robot-environment interactions [50, 30]. The effectiveness of contact-based strategies has been particularly demonstrated in robotic grasping, where accurate contact models enable reliable manipulation across diverse object geometries [4, 32].
@@ 2.2 Contact Modeling
Zhao et al[73] advanced this domain by showing that modeling contact as geometric deviations is sufficient for robust articulated object manipulation without requiring prior kinematic models. While this finding established the power of contact-based approaches, their method’s reactive, step-by-step exploration-compensation mechanism substantially limited its efficiency. Our work, TacMan-Turbo, further advances this direction by extracting local object kinematic information from these contact deviations, enabling proactive adjustments that significantly enhance manipulation efficiency while maintaining the same level of robustness.
2.3 Contact Sensing
@@
The successful implementation of TacMan-Turbo relies on accurate sensing of contacts between robots and objects, leveraging decades of advances in tactile sensing technology [65]. Based on their primary output format, tactile sensors can be categorized into two main types: force-based and geometry-based sensors, each offering distinct advantages for manipulation tasks.
@@
Force-based sensors output contact forces by converting mechanical interactions into electrical signals through various transduction mechanisms, including conductive fluid [13], resistance [64, 66], capacitance [3, 12], and others. Geometry-based sensors, conversely, directly capture the geometric characteristics of contact interfaces, with vision-based tactile sensors being prominent examples [67, 55, 27, 28, 72].
@@
Since TacMan-Turbo’s core functionality centers on contact geometry modeling for kinematic information extraction, it is naturally compatible with geometry-based sensors that directly provide the required information. Although force-based sensors can also be effectively utilized, they require an additional calibration step using Hooke’s law [18] to convert force measurements into geometric information [73]. To demonstrate TacMan-Turbo’s capabilities, in this paper, we implement our system using the GelSight sensor [67], a well-established geometry-based tactile sensing technology that provides sufficient geometric information about contact surfaces.
2.4 Contact Control
@@
Contact control, also known as tactile control, has evolved into a powerful strategy for robotic manipulation by modulating robot motion based on real-time tactile feedback. This approach was pioneered by Weiss et al[56] and subsequently advanced by Berger and Khosla [2], who provided the first practical implementation demonstrating its potential for precise interaction tasks.
@@
The fundamental challenge in tactile control lies in designing algorithms capable of effectively transitioning from current tactile observations to desired reference states. To address this challenge, researchers have developed two primary approaches: feedback control and direct mapping. Feedback control employs various controllers (e.g., PID controllers) to systematically minimize the difference between observed and desired tactile features, requiring careful selection of task-specific control parameters [26, 22, 25, 49, 35]. Direct mapping, on the contrary, transforms tactile feature differences directly into robot motion commands without explicit parameter tuning. This approach was first introduced by Chen et al[6], who proposed the concept of an inverse tactile Jacobian for 2D edge tracking using manually designed features. This foundational framework was subsequently refined and extended by numerous researchers, including Zhang et al[69] and Sikka et al[51], who demonstrated its versatility in different manipulation scenarios. More recently, Wilson et al[58] advanced the field by introducing motion primitives for cable rearrangement tasks, providing a structured approach to handle deformable objects. Zhao et al[73] made a contribution to this paradigm through their implementation of contact point registration, demonstrating robust manipulation of articulated objects despite their kinematic complexity.
@@
The latest evolution in tactile control has seen the emergence of predictive frameworks that anticipate future tactile states rather than reacting only to current observations. For example, Xu et al[63] employed learning-based techniques to predict future tactile states for reactive grasping. However, such data-driven approaches often face challenges in generalizing across diverse and dynamic tasks due to their dependence on training data distribution. Our method, TacMan-Turbo, introduces a training-free predictive control framework that takes advantage of previous sequential tactile information. By systematically extracting local kinematic information from these temporal data, our approach predicts optimal subsequent interactions and enables proactive adjustments, substantially enhancing manipulation efficiency without compromising reliability.
3 The Proactive Tactile Control
@@
The development of this framework unfolds through four interconnected components. First, Section 3.1 establishes the essential mathematical notation and preliminaries that provide the foundation for our approach. Building on this mathematical framework, Section 3.2 formally defines the articulated object manipulation problem, specifically addressing the dual optimization objectives of effectiveness and efficiency. Next, Section 3.3 introduces our novel prediction framework that analyzes sequential patterns in contact deviations to forecast articulated object states during manipulation. Finally, Section 3.4 details how these kinematic predictions are translated into optimal gripper velocity adjustments while adhering to the robot’s physical constraints and ensuring consistent contact maintenance.

### v3

@@ Abstract
Adept manipulation of articulated objects is essential for robots to operate successfully in human environments. Such manipulation requires both effectiveness—reliable operation despite uncertain object structures—and efficiency—swift execution with minimal redundant steps and smooth trajectories. Existing approaches struggle to achieve both objectives simultaneously: methods relying on predefined kinematic models lack robustness when encountering structural variations, while the tactile-informed approach achieves robust manipulation but sacrifices efficiency through reactive, step-by-step execute-and-recover cycles. To address this challenge, this paper introduces TacMan-Turbo, a proactive tactile control framework that unifies the cycles into a continuous control loop. Our key insight is to interpret tactile signals temporally in addition to spatially: instead of treating contact deviations merely as instantaneous error signals requiring immediate compensation, we analyze sequential tactile observations to reveal local kinematic information. This temporal perspective enables our controller to predict future object states and proactively modulate manipulation velocities, eliminating the previous execute-and-recover cycle. In evaluations across 200 diverse simulated articulated objects and real-world experiments, our approach maintains a near-perfect performance while significantly enhancing time efficiency, action efficiency, and trajectory smoothness (all p-values < 0.0001). These results demonstrate that tactile feedback serves a dual purpose: providing not only spatial error signals for reactive control but also critical temporal structural information that enables proactive prediction and control. By enabling robots to manipulate articulated objects both reliably and efficiently, this work advances robot capabilities for seamless operation in dynamic, human-centric environments.
Note to Practitioners
Robots operating in human-centric environments need to adeptly manipulate articulated objects (e.g., drawers and doors), despite uncertain or inaccurate kinematic models, while finishing tasks quickly and smoothly without causing damage. Traditional planners that assume known kinematics are brittle in real-world settings, and recent tactile-reactive controllers, though robust, are slow due to repeated execute-and-recover cycles. TacMan-Turbo introduces a proactive tactile control framework built on a temporal interpretation of tactile signals: by extracting local kinematic information from sequential tactile observations, the controller predicts short-horizon handle motion and proactively adjusts end-effector velocity to track these predictions, eliminating dedicated recovery phases. In evaluations across $200$ simulated articulated objects and real-world experiments, the method achieved near-perfect performance while reducing task completion time by nearly an order of magnitude compared to a strong tactile-based baseline, delivering higher action efficiency and smoother joint trajectories. Overall, TacMan-Turbo provides a training-free, drop-in upgrade for tactile manipulation of articulated objects, while preserving robustness and delivering faster, smoother, and more efficient operation without prior object kinematic models.
I Introduction
@@
Adept manipulation of articulated objects, from doors and drawers to various household items, constitutes one of the basic capabilities that robots must master to function effectively in human-centric environments [2, 3]. Success in this domain hinges on two critical objectives: effectiveness—the reliable manipulation of objects despite uncertain internal structures, and efficiency—the execution of tasks with minimal redundant movements and smooth trajectories [4, 5].
@@
Current approaches struggle to achieve both objectives simultaneously. Methods that rely on pre-provided kinematic information demonstrate efficiency in controlled scenarios but often fail when encountering discrepancies between assumed and actual object kinematics. Conversely, while the recent tactile-informed approach by Zhao et al. [1] successfully achieves robust manipulation despite kinematic uncertainty, it sacrifices efficiency—requiring several minutes to manipulate a single object due to its cautious, step-by-step exploration-compensation mechanism. This fundamental trade-off between effectiveness and efficiency remains unresolved in articulated object manipulation, as further detailed in Sec.˜II-A.
To address this challenge, we propose TacMan-Turbo, a novel proactive control framework that fundamentally reimagines how tactile feedback can guide manipulation. While leveraging Zhao et al.’s insight in using robot-object contact deviations as state representations [1], TacMan-Turbo introduces a temporal perspective: instead of treating contact deviations merely as spatial errors requiring immediate compensation, we analyze sequential tactile observations as time-series data that reveal local kinematic information. This temporal interpretation enables TacMan-Turbo to predict future object states and proactively modulate manipulation velocities to minimize predicted future deviations, streamlining the previous execute-and-recover cycle into a continuous control loop. As illustrated in LABEL:fig:teaser, this novel approach preserves the effectiveness of contact-based manipulation while enabling substantially more efficient execution.
@@
We validate TacMan-Turbo in both simulation and real-world settings. Our test scenarios systematically cover articulated objects with diverse joint types—from basic prismatic and revolute joints to complex mechanisms involving concurrent translational and rotational movements. Across $\boldsymbol{200}$ simulated articulated objects and multiple real-world experiments, TacMan-Turbo achieves the same perfect manipulation success rate as the previous tactile-informed approach while significantly improving all efficiency metrics: time efficiency, action efficiency, and trajectory smoothness (all comparisons yielding p-values < $\boldsymbol{0.0001}$). These results confirm that TacMan-Turbo successfully mitigates the long-standing effectiveness-efficiency trade-off in articulated object manipulation.
Our main contributions are summarized below:
  • A temporal interpretation of tactile signals for local kinematic inference: Rather than treating contact deviations merely as spatial errors requiring immediate compensation, we introduce a temporal perspective that analyzes sequential tactile observations as time-series data. This approach extracts local kinematic information—revealing object motion patterns from contact history—and enables prediction of future object states without requiring pre-provided kinematic models.
  • A proactive tactile control framework that unifies manipulation into a continuous process: Building on our predictive capability, we reformulate articulated object manipulation from a reactive pose-correction problem (minimizing current contact deviations) to a proactive velocity-modulation problem (minimizing predicted future deviations). This reformulation streamlines the execute-and-recover cycle into a unified process, significantly enhancing the manipulation efficiency without compromising the effectiveness.
@@
The remainder of this paper is organized as follows: Sec.˜II positions our work within the broader context of robotic manipulation research. Sec.˜III presents the theoretical framework and implementation details of TacMan-Turbo. Secs.˜IV and V evaluate TacMan-Turbo through simulation studies and real-world experiments, respectively. Sec.˜VI analyzes TacMan-Turbo’s performance characteristics and discusses future research directions. Finally, Sec.˜VII summarizes our findings and contributions.
II Related Work
@@
This section contextualizes TacMan-Turbo’s contributions within the broader landscape of robotic research. We begin by assessing current approaches to articulated object manipulation, establishing the fundamental trade-off between effectiveness and efficiency that motivates our work (Sec.˜II-A). In Sec.˜II-B, we explore contact modeling techniques, highlighting how the proposed method builds upon and advances previous work to address the effectiveness-efficiency trade-off. Next, Sec.˜II-C examines contact sensing technologies, categorizing them into force-based and geometry-based approaches while explaining their compatibility with our method. Finally, Sec.˜II-D reviews contact control strategies, tracing their evolution from early reactive feedback systems to our proposed proactive tactile control frameworks.
II-A Articulated Object Manipulation
@@
To overcome this dependency on manually provided information, subsequent research has focused on the autonomous acquisition of object kinematics. A prominent line of work leverages visual perception to infer kinematic models through geometric reasoning and motion analysis [8, 9, 10, 11, 12, 13, 14]. However, visual information alone often proves insufficient, as visually similar objects may possess fundamentally different internal mechanisms [15, 10, 1]. This inherent limitation has led researchers to integrate additional sensing modalities to resolve ambiguity. By combining multi-frame analysis with wrist-mounted force/torque sensors, more accurate parameter estimation of kinematic models becomes possible through active interaction [4, 16, 5, 17, 18, 19, 20, 21]. Nevertheless, these methods remain constrained by their underlying assumption that objects conform to predefined joint types (e.g., revolute, prismatic), limiting their effectiveness when encountering complex or unconventional mechanisms.
@@
Reinforcement learning offers an alternative paradigm that attempts to overcome the data collection bottleneck by developing manipulation strategies through repeated interactions in simulated environments [33, 34, 35, 36, 37, 38]. This approach has been significantly advanced by the creation of large-scale datasets of articulated objects [39, 40, 41, 42, 43], enabling training across diverse object categories. However, the inevitable sim-to-real gap often leads to degraded performance when these methods are deployed in real-world settings, particularly when encountering previously unseen object variations.
The limitations in both model-based and learning-based approaches recently led to the development of tactile-informed manipulation. Zhao et al. [1] introduced an approach that achieves reliable manipulation of articulated objects by compensating for robot-object contact deviations without requiring kinematic priors. While this method demonstrates robust effectiveness with unknown objects, its reactive execute-and-recover cycle results in significant time overhead. TacMan-Turbo advances this foundation through a proactive control framework that unifies the execute-and-recover cycle into a continuous control loop, significantly enhancing manipulation efficiency.
II-B Contact Modeling
Contact modeling has emerged as a fundamental approach in robotics research, providing essential information about robot-environment interactions [44, 45, 46, 47, 48]. The effectiveness of contact-based strategies has been particularly demonstrated in robotic grasping, where accurate contact models enable reliable manipulation across diverse object geometries [49, 50].
@@ II-B Contact Modeling
Zhao et al. [1] demonstrated that modeling contact as geometric deviations enables robust articulated object manipulation without requiring prior kinematic models. TacMan-Turbo advances this foundation by reinterpreting how these contact deviations are utilized: rather than treating them merely as spatial error signals requiring immediate compensation, we analyze sequential observations temporally to extract local kinematic information, enabling efficient proactive control.
II-C Contact Sensing
@@
The successful implementation of TacMan-Turbo relies on accurate sensing of contacts between robots and objects, leveraging decades of advances in tactile sensing technology [52, 53]. Based on their primary output format, tactile sensors can be categorized into two main types: force-based and geometry-based sensors, each offering distinct advantages for manipulation tasks.
@@
Force-based sensors output contact forces by converting mechanical interactions into electrical signals through various transduction mechanisms, including conductive fluid [54], resistance [55, 56, 57, 58], capacitance [59, 60, 61], and others [62, 63]. Geometry-based sensors, conversely, directly capture the geometric characteristics of contact interfaces, with vision-based tactile sensors being prominent examples [64, 65, 66, 67, 68, 69, 70, 71].
@@
Since TacMan-Turbo’s core functionality centers on contact geometry modeling for local kinematic information extraction, it is naturally compatible with geometry-based sensors that directly provide the required information. Although force-based sensors can also be effectively utilized, they require an additional calibration step using Hooke’s law [72] to convert force measurements into geometric information [1]. To demonstrate TacMan-Turbo’s capabilities, in this paper, we implement our system using the GelSight sensor [64], a well-established geometry-based tactile sensing technology that provides sufficient geometric information about contact surfaces.
II-D Contact Control
@@
Contact control, also known as tactile control, has evolved into a powerful strategy for robotic manipulation by modulating robot motion based on real-time tactile feedback. This approach was pioneered by Weiss et al. [73] and subsequently advanced by Berger and Khosla [74], who provided the first practical implementation demonstrating its potential for precise interaction tasks.
@@
The fundamental challenge in tactile control lies in designing algorithms capable of effectively transitioning from current tactile observations to desired reference states. To address this challenge, researchers have developed two primary approaches: feedback control and direct mapping. Feedback control employs various controllers (e.g., PID controllers) to systematically minimize the difference between observed and desired tactile features, requiring careful selection of task-specific control parameters [75, 76, 77, 78, 79]. Direct mapping, on the contrary, transforms tactile feature differences directly into robot motion commands without explicit parameter tuning. This approach was first introduced by Chen et al. [80], who proposed the concept of an inverse tactile Jacobian for 2D edge tracking using manually designed features. This foundational framework was subsequently refined and extended by numerous researchers, including Zhang et al. [81] and Sikka et al. [82], who demonstrated its versatility in different manipulation scenarios. More recently, Wilson et al. [83] advanced the field by introducing motion primitives for cable rearrangement tasks, providing a structured approach to handle deformable objects. Zhao et al. [1] made a contribution to this paradigm through their implementation of contact point registration, demonstrating robust manipulation of articulated objects despite their kinematic complexity.
@@
The latest evolution in tactile control has seen the emergence of predictive frameworks that anticipate future tactile states rather than reacting only to current observations. For example, Xu et al. [84] employed learning-based techniques to predict future tactile states for reactive grasping. However, such data-driven approaches often face challenges in generalizing across diverse and dynamic tasks due to their dependence on training data distribution. In contrast, TacMan-Turbo introduces a model-free, training-free predictive framework grounded in analytical interpretation of tactile sequences. Rather than learning black-box prediction models from data, we extract local kinematic information directly from sequential tactile observations through geometric analysis. This analytical approach enables TacMan-Turbo to predict future object states and make proactive adjustments without prior training.
III The Proactive Tactile Control
@@
The development of this framework unfolds through four interconnected components. First, Sec.˜III-A establishes the essential mathematical notation and preliminaries that provide the foundation for our approach. Building on this mathematical framework, Sec.˜III-B formally defines the articulated object manipulation problem, specifically addressing the dual optimization objectives of effectiveness and efficiency. Next, Sec.˜III-C introduces our novel prediction framework that analyzes sequential patterns in contact deviations to forecast articulated object states during manipulation. Finally, Sec.˜III-D details how these kinematic predictions are translated into optimal gripper velocity adjustments while adhering to the robot’s physical constraints and ensuring consistent contact maintenance.